calculate the pressure at the base of a 200-foot column with a 20 foot liquid level in the lower stripping section. the column is operating under 55 psig. specific gravity of the bottom product is 85
1
Expert's answer
2019-11-21T09:59:53-0500

The total pressure is the pressure above the fluid plus the gauge pressure of the fluid.

1 psig = 6.894 KPa.

Hence, the Column's operating pressure = 6.894×55 = 379.17 KPa

In this way, the pressure above the fluid is equal to 379.17 kPa.

The gauge pressure of the fluid is (density of fluid)x(depth in fluid)x(acceleration due to gravity:g).

The specific gravity is equal to the ratio of the density of the fluid to that of water.

In our case, the density of the fluid is equal to 0.85×103 kg/m3.

In this way, the gauge pressur is equal to 20ft × 30.48 cm×0.85×103 kg/m3×9.8=50.73 kPa

The total pressure is equal to 379.17 KPa+50.73 KPa=430 kPa


Answer

430 kPa